# Quantum Gravity

### v1

@@ 5.1 Bubble diagram
It is convenient to split the calculation in two steps:
@@
Since the loop momentum $p$ is Euclidean, the limit $\lambda^{\prime}\rightarrow+\infty$ can be evaluated immediately by means of (2.8), (2.17) or (3.2). We obtain
@@
We overcome the difficulty by adding and subtracting integration paths. Consider the $p_{4}$ integral. Its integration domain is the line $C$ (see fig. 2 – left side). We can rewrite the integral on $C$ as the sum of the integral on the closed curve $\gamma$, plus the integral on the line $C^{\prime}$ (fig. 2 – right side). The segments at infinity do not contribute, since they are located inside the cone $\mathcal{C}^{\prime}$ where $\lambda^{-1}Q(\lambda P(p-k_{\text{M}}))$ converges to $P(p-k_{\text{M}})$, as follows from formula (2.12) [^5] Recall that we are using the dimensional regularization. This means that when an integral is ultraviolet divergent, we split it into a sum of integrals that admit convergence domains in the complex plane of the dimension $D$ [23]. Then the arguments are applied to each of them separately..
The integrand of (5.4) has poles at $p^{2}+m^{2}=0$ in the complex $p_{4}$ plane. One of them may fall inside the curve $\gamma$. We need to distinguish the case where this happens from the opposite case. The integral on $C^{\prime}$, instead, is centered on the Euclidean region, so we can take the limit $\lambda\rightarrow+\infty$ straightforwardly on it by means of (2.17).
@@
Let us begin by evaluating $(a)$ and $(b)$ in the simplified case $\mathbf{k}=m=0$. The residue theorem gives [^6] Note that the integral is in $p_{4}=-ip^{0}$, not $p^{0}$, so there is an extra factor $-i$.
@@
Since the argument of $Q$ is real, we can take the limit $\lambda\rightarrow+\infty$ by means of formula (2.17). We obtain
@@
which is precisely the result obtained in the case of PVPs (apart from the factor $V$), as given in formula (A.2).
@@
What is important is that the argument of $Q$ remains real, so we can still use formula (2.17) to evaluate the limit $\lambda\rightarrow+\infty$. The result is
@@
The singularity due to the first term inside the parentheses is regulated by the principal value inherited from (2.17). Instead, the second term is nonsingular for $\omega_{\mathbf{p}}\leqslant k^{0}$.
@@
which is the same as in the case of PVPs, formula (A.2).
If one particle is physical, like $\varphi$ in (2.3), and the other one is $\phi$ with the second option of (2.5) for $Q$, the result does not change. The simplest parametrization of the circulating momenta gives (5.4) directly, whence the rest follows as before. If we start from the parametrization
@@
we have to note that the integration on $p_{4}$ must be deformed to complex values in order to leave the right pole on one side and the left pole on the other side. Then the argument of $Q$ is not everywhere real, so the limit $\lambda\rightarrow+\infty$ cannot be taken directly. On the other hand, one can easily see that (5.17) is equivalent to (5.4), because the difference is a closed path that encircles no pole.
5.2 Triangle diagram
@@
We introduce the parameter $\lambda$ and split the local limit $\lambda\rightarrow+\infty$ in three steps. First, we take it inside $G_{\text{nl}}(p)$, then in $G_{\text{nl}}(p-k_{\text{M}})$ and finally in $G_{\text{nl}}(p-q_{\text{M}})$. The calculation will tell us to what extent it is legitimate to do so.
@@
Now we use the residue theorem to integrate the loop energy $p^{4}=-ip^{0}$ along the closed curve $\gamma$ made of the lines $p^{0}=0$ and $p^{0}=k^{0}$, plus segments at infinity. On the pole $p^{0}=\omega_{\mathbf{p}}$, which contributes for $\omega_{\mathbf{p}}<k^{0}$, the arguments of the functions $Q$ are real, so we can use formula (2.17). We then obtain the first contribution, which is
@@
$\mathcal{P}$ denotes the Cauchy principal value, and the subscripts $a,b,\ldots$ range over the values $1$, $2$ and $3$, while $\{e_{a}\}=\{0,-k^{0},-q^{0}\}$, $\{\omega_{a}\}=\{\omega_{\mathbf{p}},\omega_{\mathbf{p-k}},\omega_{\mathbf{p-q}}\}$.
@@
Using the residue theorem once more on $(b_{1})$, we obtain an integral on $\mathbf{p}$ that we do not report here. We just remark that, before taking the limit $\lambda\rightarrow\infty$, its integrand is regular. Specifically, the residues at the poles mutually compensate for every $\lambda<\infty$. This means that we can adopt the prescription we want for those poles. We choose the principal value. Then, we take the limit $\lambda\rightarrow\infty$ by means of (2.17), noting that the arguments of $Q$ are real on the poles. We find that the contribution of $\gamma^{\prime}$ is
@@
Now the limit $\lambda\rightarrow+\infty$ acts on $\lambda/Q(\lambda P(p))$. Nevertheless, we cannot evaluate it directly by means of (2.17), because if we do so, we find an unprescribed expression:
@@
The problem just outlined indicates that we have jumped to (5.21) too quickly, since the integral of (5.20) is well defined before taking the limit. To avoid the trouble, we can evaluate (5.20) as follows. First, we move the external energies slightly away from the real axis: $k^{0}\rightarrow k^{0}+i\epsilon$, $q^{0}\rightarrow q^{0}+i\epsilon^{\prime}$, with $\epsilon$ and $\epsilon^{\prime}$ real (not necessarily positive) and small. Then we use (2.17) to evaluate the limit $\lambda\rightarrow+\infty$. When we apply the residue theorem, we find expressions like (5.23) with $x\rightarrow x-i\epsilon^{\prime}$ and $y\rightarrow y+i\epsilon$, which do vanish.
@@
$\mathcal{T}_{\text{M}}^{\text{loc}}(k_{\text{M}},q_{\text{M}})$ $=$ $(a)+(b)=-i\int\frac{\mathrm{d}^{D-1}\mathbf{p}}{(2\pi)^{D-1}}\frac{\mathcal{P}^{12}\mathcal{P}^{13}+\mathcal{P}^{21}\mathcal{P}^{31}+\text{cycl}}{8\omega_{\mathbf{p}}\omega_{\mathbf{p-k}}\omega_{\mathbf{p-q}}}$ (5.25)
$+i\pi^{2}\int\frac{\mathrm{d}^{D-1}\mathbf{p}}{(2\pi)^{D-1}}\frac{\delta(q^{0}-\omega_{\mathbf{p}}-\omega_{\mathbf{p-q}})\delta(q^{0}-k^{0}-\omega_{\mathbf{p-k}}-\omega_{\mathbf{p-q}})}{8\omega_{\mathbf{p}}\omega_{\mathbf{p-k}}\omega_{\mathbf{p-q}}}.$
@@ 6.1 PVP-QG
$S_{\text{QG}}(\tilde{g},\phi,\chi,\Phi)=\tilde{S}_{\text{HE}}(\tilde{g})+S_{\phi}(\tilde{g}+\psi,\phi)+S_{\chi}(\tilde{g},\chi)+S_{m}(\tilde{g}\mathrm{e}^{\kappa_{\phi}\phi}+\psi\mathrm{e}^{\kappa_{\phi}\phi},\Phi),$ (6.2)
@@
Expanding around the flat-space metric $\eta_{\mu\nu}=$diag(1,1,1,1)$1$(1,-1,-1,-1)( 1 , - 1 , - 1 , - 1 ), the $\chi_{\mu\nu}$ propagator of (6.2) is
@@
where and the subscript “PVP” is there to remind us that $\chi_{\mu\nu}$ must be treated as a PVP inside diagrams.
The projection applies at the classical level as well. The true, classical theory is neither (6.1) nor (6.2). It is obtained by collecting the tree diagrams with no $\chi_{\mu\nu}$ external legs [19]. The result is
@@ 6.2 Kinetic Lagrangians of Proca and Pauli-Fierz AL-QFTs
As anticipated, the new “propagator” (6.9) has no pole, hence it does not actually propagate degrees of freedom. We can view it as the propagator of a NL-PVP. In the local limit, the function $\sigma^{2}$ tends to one, so (6.9) tends to (6.8) (at the tree level), by formula (2.17).
@@
$+2\chi^{\mu\nu}D_{\mu}\tilde{Q}(D_{\rho}D^{\rho}+m_{\chi}^{2})D_{\nu}\chi-2\chi^{\rho\nu}D_{\mu}\tilde{Q}(D_{\rho}D^{\rho}+m_{\chi}^{2})D_{\rho}\chi_{\nu}^{\mu}$
@@ 6.4 Vertices
Now we show that the vertices obtained by expanding (6.12) around flat space are well defined (see [28] and [6]). Write
@@
$f(x,y)=\frac{f(x)-f(y)}{x-y}$
denotes the incremental ratio of the function $f$, and the arrow on $\overleftarrow{\Box}$ means that the box acts to the very left, beyond $V_{\mu}\partial^{\mu}+W$. In momentum space, the ratio is calculated with respect to the right and left momenta.
Since the incremental ratio of an entire function is an entire function (of two variables), no singularity appears. Moreover, with $f=Q$ the ratio converges to
$\frac{Q(-p^{2}+m_{\chi}^{2})-Q(-q^{2}+m_{\chi}^{2})}{-p^{2}+q^{2}}\rightarrow\mathcal{P}\frac{\frac{1}{p^{2}-m_{\chi}^{2}}-\frac{1}{q^{2}-m_{\chi}^{2}}}{p^{2}-q^{2}}=-\mathcal{P}\frac{1}{p^{2}-m_{\chi}^{2}}\frac{1}{q^{2}-m_{\chi}^{2}}$ (6.14)
@@
Formula (6.15) shows that the second-order vertex is the incremental ratio of the incremental ratio, keeping one variable fixed (which one being immaterial). In particular, the vertex is well defined and tends to its local limit fast enough. One can proceed similarly for the rest of the expansion.
To derive the loop integrals of the theory (6.12), we proceed as follows. First, we integrate out $\chi_{\mu\nu}$. This generates “functional diagrams” built by means of the covariantized versions of the propagators (6.9), where $Q$ and $\sigma$ are functions of $D_{\mu}D^{\mu}+m^{2}$. Later, we expand the metric tensor $g_{\mu\nu}$ around flat space. The expansion also acts inside the covariantized $\chi_{\mu\nu}$ propagators, and is treated by means of identities like (6.13). Once this is done, we arrive at the loop integrals, which are well defined.
@@ 6.5 Arbitrariness and renormalization
The nonlocal theories of the literature [5, 7, 8] are super-renormalizable. Unfortunately, their renormalization properties do not extend to (6.1), which is strictly renormalizable [6]. In view of this, AL-QG is not expected to be renormalizable either. For the reasons that we explain below, we do not think that this is a serious liability.
Although PVP-QG is unique [13], the nonlocal extension PVP-QG $\rightarrow$ AL-QG is not. Any entire function like those considered in the literature [5, 7, 8] can be used for $h$ inside the second option for the function $Q$ of (2.5). This leads to an infinite arbitrariness. The arbitrariness likely turned on by renormalization, together with the one associated with the adjustments of the multi-threshold contributions mentioned in the previous section, is not worse than that.
@@
The main weakness of nonlocal quantum field theory is the lack of a fundamental principle for selecting the form factors that modify the propagators. In our approach this problem is addressed by the very existence of the local limit in Minkowski spacetime. If we view that as the missing guiding principle, the candidate theories of the universe are the local limits themselves.
The arbitrariness of AL-QFT is reminiscent of the one of off-the-mass shell physics [18]. In the latter, the extra parameters are not properties of the fundamental interactions, but describe the environment where the phenomenon is observed, such as the experimental apparatus and the observer itself. We could say that they describe the quantum/classical interplay between the phenomenon and the rest of the universe. It would be interesting to uncover the map relating the arbitrariness of AL-QFT to the one of the off-shell approaches to QFT of ref.s [18]. We postpone this task, because it is beyond the scope of this paper.
We stress that the non uniqueness of the nonlocal deformation does not necessarily imply a lack of predictivity. The number of parameters impacting the phenomenon we want to observe is hopefully finite. Once they are identified, they can be fixed by sacrificing an equal number of initial measurements, after which every other measurement is predicted efficiently.

### v3

@@ 4.4 Local limit: bubble diagram
As recalled in appendix B, PVPs are defined by a thoroughlly new diagrammatics – i.e., not just the usual diagrams with propagators replaced by the principal value appearing in (4.7). Because of this, before concluding that the local limit with the second choice of (2.6) gives a theory with PVPs, we must prove that the loop diagrams somehow “know what to do” by themselves. That is to say, for some mysterious reason they are already equipped with the instructions to implement the diagrammatic rules of [10, 11, 12], instead of those that lead to (B.3). Luckily, this turns out to be true [^4] Besides, we already know that they cannot give (B.3), because unitarity rules out local limits with ghosts..
It is convenient to split the calculation of the bubble diagram in two steps:
@@
Since the loop momentum $p$ is Euclidean, the limit $\lambda^{\prime}\rightarrow+\infty$ can be evaluated immediately by means of (A.3), (A.12) or (3.2). We obtain
@@
We overcome the difficulty by adding and subtracting integration paths. Consider the $p_{4}$ integral. Its integration domain is the line $C$ (see fig. 1 – left side). We can rewrite the integral on $C$ as the sum of the integral on the closed curve $\gamma$, plus the integral on the line $C^{\prime}$ (fig. 1 – right side). The segments at infinity do not contribute, since they are located inside the cone $\mathcal{C}^{\prime}$ where $\lambda^{-1}Q(\lambda P(p-k_{\text{M}}))$ converges to $P(p-k_{\text{M}})$, as follows from formula (A.7) [^5] Recall that we are using the dimensional regularization. This means that when an integral is ultraviolet divergent, we split it into a sum of integrals that admit convergence domains in the complex plane of the dimension $D$ [22]. Then the arguments are applied to each of them separately..
The integrand of (4.8) has poles at $p^{2}+m^{2}=0$ in the complex $p_{4}$ plane. One of them may fall inside the curve $\gamma$. We need to distinguish the case where this happens from the opposite case. The integral on $C^{\prime}$, instead, is centered on the Euclidean region, so we can take the limit $\lambda\rightarrow+\infty$ straightforwardly on it by means of (A.12).
@@
Let us begin by evaluating $(a)$ and $(b)$ in the simplified case $\mathbf{k}=m=0$. The residue theorem gives [^6] Note that the integral is in $p_{4}=-ip^{0}$, not $p^{0}$, so there is an extra factor i-i.
@@
Since the argument of $Q$ is real, we can take the limit $\lambda\rightarrow+\infty$ by means of formula (A.12). We obtain
@@
which is precisely the result obtained in the case of PVPs (apart from the factor $V$), as given in formula (B.2).
@@
What is important is that the argument of $Q$ remains real, so we can still use formula (A.12) to evaluate the limit $\lambda\rightarrow+\infty$. The result is
@@
The singularity due to the first term inside the parentheses is regulated by the principal value inherited from (A.12). Instead, the second term is nonsingular for $\omega_{\mathbf{p}}\leqslant k^{0}$.
@@
which is the same as in the case of PVPs, formula (B.2).
If one particle circulating in the loop is physical, like $\varphi$ in (2.4), and the other one is $\phi$ with the second option of (2.6) for $Q$, the result does not change. The simplest parametrization of the circulating momenta gives (4.8) directly, whence the rest follows as before. If we start from the parametrization
@@
we have to note that the integration on $p_{4}$ must include complex values in order to leave the right pole on one side and the left pole on the other side. Then the argument of $Q$ is not everywhere real, so the limit $\lambda\rightarrow+\infty$ cannot be taken directly. On the other hand, one can easily see that (4.21) is equivalent to (4.8), because the difference is a closed path that encircles no pole.
4.5 Local limit: triangle diagram
@@
We introduce the parameter $\lambda$ and split the local limit $\lambda\rightarrow+\infty$ in three steps. First, we take the limit inside $G_{\text{nl}}(p)$, then in $G_{\text{nl}}(p-k_{\text{M}})$ and finally in $G_{\text{nl}}(p-q_{\text{M}})$. The calculation will tell us to what extent it is legitimate to do so.
@@
Now we use the residue theorem to integrate the loop energy $p^{4}=-ip^{0}$ along the closed curve $\gamma$ made of the lines $p^{0}=0$ and $p^{0}=k^{0}$, plus segments at infinity. On the pole $p^{0}=\omega_{\mathbf{p}}$, which contributes for $\omega_{\mathbf{p}}<k^{0}$, the arguments of the functions $Q$ are real, so we can use formula (A.12). We then obtain the first contribution, which is
@@
and the subscripts $a,b,\ldots$ range over the values $1$, 22 and $3$, while $\{e_{a}\}=\{0,-k^{0},-q^{0}\}$, $\{\omega_{a}\}=\{\omega_{\mathbf{p}},\omega_{\mathbf{p-k}},\omega_{\mathbf{p-q}}\}$.
@@
Using the residue theorem once more on $(b_{1})$, we obtain an integral on $\mathbf{p}$ that we do not report here. We just remark that, before taking the limit $\lambda\rightarrow\infty$, its integrand is regular. Specifically, the residues at the poles mutually compensate for every $\lambda<\infty$. This means that we can adopt the prescription we want for those poles. We choose the principal value. Then, we take the limit $\lambda\rightarrow\infty$ by means of (A.12), noting that the arguments of $Q$ are real on the poles. We find that the contribution of $\gamma^{\prime}$ is
@@
Now the limit $\lambda\rightarrow+\infty$ acts on $\lambda/Q(\lambda P(p))$. Nevertheless, we cannot evaluate it directly by means of (A.12), because if we do so, we find an unprescribed expression:
@@
The problem just outlined indicates that we have jumped to (4.25) too quickly, since the integral of (4.24) is well defined before taking the limit. To avoid the trouble, we can evaluate (4.24) as follows. First, we move the external energies slightly away from the real axis: $k^{0}\rightarrow k^{0}+i\epsilon$, $q^{0}\rightarrow q^{0}+i\epsilon^{\prime}$, with $\epsilon$ and $\epsilon^{\prime}$ real (not necessarily positive) and small. Then we use (A.12) to evaluate the limit $\lambda\rightarrow+\infty$. When we apply the residue theorem, we find expressions like (4.27) with $x\rightarrow x-i\epsilon^{\prime}$ and $y\rightarrow y+i\epsilon$, which do vanish.
@@
$\mathcal{T}_{\text{M}}^{\text{loc}}(k_{\text{M}},q_{\text{M}})$ $=$ $-i\int\frac{\mathrm{d}^{D-1}\mathbf{p}}{(2\pi)^{D-1}}V\frac{\mathcal{P}^{12}\mathcal{P}^{13}+\mathcal{P}^{21}\mathcal{P}^{31}+\text{cycl}}{8\omega_{\mathbf{p}}\omega_{\mathbf{p-k}}\omega_{\mathbf{p-q}}}$ (4.29)
$+i\pi^{2}\int\frac{\mathrm{d}^{D-1}\mathbf{p}}{(2\pi)^{D-1}}V\frac{\delta(q^{0}-\omega_{\mathbf{p}}-\omega_{\mathbf{p-q}})\delta(q^{0}-k^{0}-\omega_{\mathbf{p-k}}-\omega_{\mathbf{p-q}})}{8\omega_{\mathbf{p}}\omega_{\mathbf{p-k}}\omega_{\mathbf{p-q}}}.\quad$
@@ 5.1 PVP-QG
$S_{\text{QG}}(\tilde{g},\phi,\chi,\Phi)=\tilde{S}_{\text{HE}}(\tilde{g})+S_{\phi}(\tilde{g}+\psi,\phi)+S_{\chi}(\tilde{g},\chi)+S_{m}(\tilde{g}\mathrm{e}^{\kappa_{\phi}\phi}+\psi\mathrm{e}^{\kappa_{\phi}\phi},\Phi).$ (5.2)
@@
Expanding around the flat-space metric $\eta_{\mu\nu}=$diag(1,1,1,1)(1,-1,-1,-1), the $\chi_{\mu\nu}$ propagator of (5.2) is
@@
where the subscript “PVP” is there to remind us that $\chi_{\mu\nu}$ must be treated as a PVP inside diagrams.
The projection applies at the classical level as well. The true, classical theory is neither (5.1) nor (5.2). It is obtained by collecting the tree diagrams with no $\chi_{\mu\nu}$ external legs and only $\chi_{\mu\nu}$ internal legs [18]. The result is
@@ 5.2 Kinetic Lagrangians of Proca and Pauli-Fierz AL-QFTs
As anticipated, the new “propagator” (5.9) has no pole, hence it does not actually propagate degrees of freedom. We can view it as the propagator of a NL-PVP. In the local limit, the function $\sigma^{2}$ tends to one, so (5.9) tends to (5.8) (at the tree level), by formula (A.12).
@@
$+2\chi^{\mu\nu}D_{\mu}\tilde{Q}(D_{\rho}D^{\rho}+m_{\chi}^{2})D_{\nu}\chi-2\chi^{\sigma\nu}D_{\mu}\tilde{Q}(D_{\rho}D^{\rho}+m_{\chi}^{2})D_{\sigma}\chi_{\nu}^{\mu}$
@@ 5.4 Vertices
Now we show that the vertices obtained by expanding (5.12) around flat space are entire functions (see [25] and [6]). Write
@@
$f(x,y)\equiv\frac{f(x)-f(y)}{x-y}$ (5.14)
[…]
$\frac{Q^{-1}(-p^{2}+m_{\chi}^{2})-Q^{-1}(-q^{2}+m_{\chi}^{2})}{-p^{2}+q^{2}}\rightarrow\mathcal{P}\frac{\frac{1}{p^{2}-m_{\chi}^{2}}-\frac{1}{q^{2}-m_{\chi}^{2}}}{p^{2}-q^{2}}=-\mathcal{P}\frac{1}{p^{2}-m_{\chi}^{2}}\frac{1}{q^{2}-m_{\chi}^{2}}$ (5.15)
@@
Formula (5.16) shows that the second-order vertex is the incremental ratio of the incremental ratio,
$f(x,y,z)=\frac{f(x,z)-f(y,z)}{x-y},$ (5.17)
keeping one variable fixed (which one being immaterial). In particular, the vertex is well defined and tends to its local limit fast enough. One can proceed similarly for the rest of the expansion.
To derive the loop integrals of the theory (5.12), we proceed as follows. First, we integrate out $\chi_{\mu\nu}$. This generates “functional diagrams” built by means of the covariantized versions of the propagators (5.9), where $1/Q$ and $\sigma$ are (entire) functions of $D_{\mu}D^{\mu}+m^{2}$. Later, we expand the metric tensor $g_{\mu\nu}$ around flat space. The expansion also acts inside the covariantized $\chi_{\mu\nu}$ propagators, and is treated by means of identities like (5.13). By the results of this section, it generates only entire functions, through repeated incremental ratios. Once the expansion is worked out to the order of our interest, we are ready to study the loop integrals. They are well defined, because the vertices are entire functions, the propagators of physical particles are standard ones and the propagators of PVPs are also entire functions.
Summarizing,
  1. 1.
    the covariantized Lagrangians of the theories considered here generate entire functions $1/Q$ and $\sigma$ of the covariant box, upon integrating out the nonlocal purely virtual fields;
  2. 2.
    the vertices, obtained by expanding the covariant derivatives as simple derivatives plus corrections proportional to fields, involve repeated incremental ratios;
  3. 3.
    the repeated incremental ratios of entire functions, formulas (5.14) and (5.17), are entire functions of more variables, because they can be expanded in power series of all their variables at the same time [^9] Their expressions show that the zeros of the denominators are compensated by the zeros of the numerators, so no singularity occurs.;
  4. 4.
    because of this, the diagrammatic structure is the one exhibited in formula (2.9).
That is to say, the asymptotically local theories we have formulated, such as the nonlocal deformation (5.12) of quantum gravity with purely virtual particles, lead to Feynman diagrams where the vertices and the propagators are made of entire functions, apart from the propagators of the physical particles, which have the usual poles.
@@ 5.5 Arbitrariness and renormalization
The nonlocal theories of the literature [5, 7, 8] are super-renormalizable. Unfortunately, their renormalization properties do not extend to nonlocal deformations of strictly renormalizable theories, such as (5.1) [6]. In view of this, AL-QG is not expected to be renormalizable either. For the reasons that we explain below, we do not think that this is a serious liability.
Although PVP-QG is unique [13], the nonlocal extension PVP-QG $\rightarrow$ AL-QG is not. Any entire function like those considered in the literature [5, 7, 8] can be used for $h$ inside the second option for the function $Q$ in (2.6). This leads to an infinite arbitrariness. The arbitrariness likely turned on by renormalization, together with the one associated with the adjustments of the multi-threshold contributions mentioned in the previous section, is not worse than that.
@@
The main weakness of nonlocal quantum field theory is the lack of a fundamental principle for selecting the form factors that modify the propagators. In our approach this problem is addressed by the very existence of the local limit in Minkowski spacetime, in the sense that we can view that as the missing guiding principle. Then the candidate theories of the universe are the local limits themselves.
The arbitrariness of AL-QFT is reminiscent of the one of off-the-mass shell physics [20]. In the latter, the extra parameters are not properties of the fundamental interactions, but describe the environment where the phenomenon is observed, such as the experimental apparatus and the observer itself. We could say that they account for the quantum/classical interplay between the phenomenon and the rest of the universe. It would be interesting to uncover the map relating the arbitrariness of AL-QFT to the one of the off-shell approaches to QFT of ref.s [20]. We postpone this task, because it is beyond the scope of this paper.
We stress that the non uniqueness of the nonlocal deformation does not imply a lack of predictivity. The number of parameters impacting the phenomenon we want to observe is hopefully finite. Once they are identified, they can be fixed by sacrificing an equal number of initial measurements, after which every other measurement is predicted efficiently.